# Whirling injections, surjections, and other functions between finite sets

### v3

@@ Abstract.
This paper analyzes a certain action called “whirling” that can be defined on any family of functions between two finite sets equipped with a linear (or cyclic) ordering. Many maps of interest in dynamical algebraic combinatorics, such as rowmotion of order ideals, can be represented as a composition of “toggling” involutions, each of which modifies its object only locally. Similarly whirling is made up of locally-acting whirling maps which directly generalize toggles, but cycle through more than two possible outputs. In this first paper on whirling, we consider it as a map on subfamilies of functions between finite sets.
@@
Keywords: dynamical algebraic combinatorics, homomesy, injection, noncrossing partition, RG-word, set partition, surjection, toggles, toggle group, whirling, whirls
@@ 1. Introduction
In this paper, we explore the orbits associated with an action that can be defined on any family $\mathcal{F}$ of functions $f$ from one finite set $S$ to another finite set $T$, where $S$ comes with a linear ordering and $T$ comes with a cyclic ordering. Without loss of generality we may take $S=[n]:=\{1,2,\dots,n\}$ with its usual ordering and $T=[k]$ with the mod $k$ cyclic ordering. The action is generated by an operation $\mathbf{w}:\mathcal{F}\rightarrow\mathcal{F}$ (“the whirling map”) that is in turn defined as the composition of simpler maps $w_{i}:\mathcal{F}\rightarrow\mathcal{F}$ (“whirling at $i$”, with $i$ in $[n]$) that repeatedly add $1\bmod k$ to the value of $f$ at $i$ until we get a function in $\mathcal{F}$. Our main results are examples of the homomesy [^1] Greek for “same middle” phenomenon, introduced by the second and third authors [PR15] and defined as follows.
@@
Although introduced relatively recently, the homomesy phenomenon has been discovered to be widespread throughout combinatorial dynamical systems. See [Rob16] for a survey article. Since it is a new area of research, there is still much to learn about it, including the best techniques for proving homomesy, what types of actions and statistics commonly yield homomesy, and what consequences can arise from homomesy.
Early on, almost all proven homomesy results were for actions whose order was straightforwardly computable in general and relatively small compared to the size of the ground set $\mathcal{S}$. So it is rarer and notable to find homomesies for actions with unpredictable orbit sizes, as is the case with the maps in this paper, which also generally have unknown order. Other such examples can be found in [EFG+16, ELM+23]. In some cases, a statistic which is always an integer has a non-integer average across every orbit, which means all of the orbit sizes must be multiples of the constant average’s denominator. This consequence of homomesy appears in Corollary 2.22.
The whirl maps $w_{i}$ are a generalization of the toggle maps that are heavily studied in dynamical algebraic combinatorics. Toggles were first introduced as a way of writing the rowmotion map as a composition of simpler operations on order ideals [CF95]. Striker later generalized them to a much wider variety of settings [Str18]. There have been many interesting homomesy and periodicity phenomena found for actions defined in terms of toggles; see e.g., [PR15, Rob16, Had21, EFG+16, JR18, ELM+23]. In recent work, Plante and the third author consider a whirling action on the set of $k$-bounded $P$-partitions of a poset $P$. They prove this to be in equivariant bijection with rowmotion acting on order ideals of the poset $P\times[k]$ for any finite poset $P$. They then leverage this to obtain periodicity and homomesy results for rowmotion acting on the poset ${\sf V}\times[k]$, which has surprisingly good dynamical properties [PR24]. A key point here is that beyond the intrinsic interest of whirling as a generalization of toggling, its study can facilitate our understanding of rowmotion questions that arise independently.
In Section 2 we consider the whirling action on injections and on surjections between two finite sets. In both cases, we prove that within each orbit, any two elements $1,2,\dots,k$ of the codomain appear as outputs of functions the same number of times. This was originally conjectured by the second author. We also prove this for a family of functions that naturally generalizes injections and conjecture it for a similar family that generalizes surjections. Then in Section 3, we explore the whirling action on various families of restricted growth words, which are functions that encode set partitions.
@@ 1.1. Acknowledgements
We have benefitted from useful discussions with Yue Cai, David Einstein, Max Glick, Shahrzad Haddadan, Sam Hopkins, Thomas McConville, Matthew Plante, Elizabeth Sheridan Rossi, Richard Stanley, Jessica Striker, and Nathan Williams. Computations leading to the initial conjectures were done in Sage [S+16].
@@ 2.1. The whirling map on m-injections and m-surjections
A function $f:S\rightarrow T$ is $m$-injective if every element $t\in T$ appears as an output of $f$ at most $m$ times, i.e., #$\#f^{-1}(t)\leq m$ for every $t\in T$. An $m$-injective function is also called an $m$-injection. Injections are the same as 1-injections.
@@
A function $f:S\rightarrow T$ is $m$-surjective if every element $t\in T$ appears as an output of $f$ at least $m$ times, i.e., #$\#f^{-1}(t)\geq m$ for every $t\in T$. An $m$-surjective function is also called an $m$-surjection. Surjections are the same as 1-surjections.
@@
For $f\in\mathcal{F}\subseteq[k]^{[n]}$ define a map $w_{i}:\mathcal{F}\rightarrow\mathcal{F}$, called whirling at index $i$, as follows: repeatedly add 1 (mod $k$) to the value of $f(i)$ until we get a function in $\mathcal{F}$. The new function is $w_{i}(f)$. (Here we represent our residues (mod $k$) within $[k]$ rather than $\{0,1,2,\dotsc,k-1\}$.)
@@
Let $\mathcal{F}=\operatorname{Inj}_{2}(6,4)$ and $f=422343$. To compute $w_{3}(f)$, we first add 1 to $f(3)=2$ and get the function $423343$, which is not 2-injective since 3 appears as an output three times. So, we add 1 again to the third position and get $424343$, which is also not 2-injective. Adding 1 (mod 4) once more gives a 2-injective function $421343=w_{3}(f)$.
@@
Remark 2.6.
In the case that $\mathcal{F}=\operatorname{Sur}_{m}(n,k)$, the only possible reason we could not add 1 mod $k$ to $f(i)$ is because the value $f(i)$ is only an output of the function $m$ times. (Note that in the previous sentence, the first $f(i)$ is a dynamic value, subject to change as we update $f$, while the second is a value, associated with a particular $f$; we trust that this dual usage will not create confusion.) Therefore, in this case, $w_{i}$ either adds 1 mod $k$ to $f(i)$, or it leaves $f$ alone. This is not the case when $\mathcal{F}=\operatorname{Inj}_{m}(n,k)$.
@@
$41652753\stackrel{{\scriptstyle w_{1}}}{{\longmapsto}}4753\stackrel{{\scriptstyle w_{2}}}{{\longmapsto}}4153\stackrel{{\scriptstyle w_{3}}}{{\longmapsto}}4163\stackrel{{\scriptstyle w_{4}}}{{\longmapsto}}$ 4165
@@
On any family $\mathcal{F}$, the map $w_{i}$ (and thus $\mathbf{w}$ also) is invertible. Given $f\in\mathcal{F}$, we get $w_{i}^{-1}(f)$ by repeatedly subtracting 1 (mod $k$) from the value of $f(i)$ until we get a function in $\mathcal{F}$.
@@
Note that Conjecture 2.12 holds for $k=2$ by Theorem 2.11 since $\operatorname{Sur}_{m}(n,2)=\operatorname{Inj}_{n-m}(n,2)$. The conjecture has been verified by a computer for all triples $(m,n,k)$ where $n\leq 10$ (for which there are finitely many relevant triples since $\operatorname{Sur}_{m}(n,k)=\O$ if $mk>n$).
2.2. The proof of Theorem 2.11 for injections
In this subsection, we prove Theorem 2.11 when $\mathcal{F}=\operatorname{Inj}_{1}(n,k)$. Even though this is a special case of $m$-injections, it is easier to understand the proof in this simpler situation. We will utilize much of the same notation and terminology for the other cases.
Let $\mathcal{O}$ be an orbit under the action of $\mathbf{w}$ on $\mathcal{F}$. We draw a board for the orbit $\mathcal{O}$ by placing some $f\in\mathcal{O}$ on the top line. The function in row $i+1$ is $\mathbf{w}^{i}(f)$ for $i\in[0,\ell(\mathcal{O})-1]$, where $\ell(\mathcal{O})$ is the length of $\mathcal{O}$. For example, in Figure 1, we show a board for an orbit on $\mathcal{F}=\operatorname{Inj}_{1}(3,6)$. From this board, we see that if we let $f=621$, then $\mathbf{w}(f)=342$, $\mathbf{w}^{2}(f)=563$, and so on, and that $\mathbf{w}^{10}(f)=f$.
Notice that within an orbit board, if $f$ is a given line, then the “partially whirled element” $(w_{i}\circ\cdots\circ w_{1})(f)$ is given by the first $i$ numbers on the line below $f$ and the last $n-i$ numbers of $f$. (We say the line “below” the bottom line is the top line, as we consider the orbit board to be cylindrical.) For instance, in the first two lines of Figure 1, $f=621$, $w_{1}(f)=321$, and $(w_{2}\circ w_{1})(f)=341$.
@@
6 2 1
3 4 2
5 6 3
1 2 4
3 5 6
4 1 2
5 3 4
6 5 1
2 6 3
4 1 5
Figure 1. Left: The orbit under the action of $\mathbf{w}$ on $\operatorname{Inj}_{1}(3,6)$ containing $f=621$. Right: This same orbit but with colors representing the $[6]$-chunks, defined below.
We use the term reading the orbit board to refer to this action where we start at a certain position $P$ of the board, and continue to the right until we reach the end of the line, and then go to the leftmost position of the line below and continue. This is because it is exactly like reading a book (except that continuing past the bottom line means returning to the top line). When we refer to a certain position being $x$ positions “before” or “after” the position $P$, or say the “previous” or “next” position, we always mean in the reading order.
@@
[…]
@@
Consider the following orbit on $\operatorname{Inj}_{1}(4,5)$. Let $P$ be the position in the second row and second column, shown below surrounded by a black rectangle. Then $(P,[1,4])$ consists of the four positions circled in red. Also $(P,[0,4])$ is $(P,[1,4])$ together with $P$, while $(P,[-1,2])$ is the second row. As the orbit is cylindrical, the bottom right corner is both $(P,14)$ and $(P,-6)$. Note that $P$ refers only to the position, not the value in that position. So $P\not=(P,5)$ since they are different positions, even though both contain the value 3. Similarly, we will never write, e.g., $P=3$.
@@
Note that if $P$ is in row $i$ and column $j$, then $(P,[1,n])$ always consists of all positions to the right of $P$ in row $i$, followed with the leftmost $j$ positions of row $i+1$. Also note that $P([-n,-1])$ consists of the rightmost $n-j+1$ positions of row $i-1$ followed by all $j-1$ positions left of $P$ in row $i$.
@@
    There is exactly one occurrence of $i+1\bmod k$ within $(P,[1,n])$.
@@
    There is exactly one occurrence of $i-1\bmod k$ within $(P,[-n,-1])$.
@@
To prove (1), suppose position $P$ is in column $j$. So $f(j)=i$ for some $f\in\mathcal{O}$. Then $(P,[1,n])$ contains the multiset of outputs of $(w_{j}\circ\cdots\circ w_{1})(f)$ in some order. If $(w_{j-1}\circ\cdots\circ w_{1})(f)$ does not have $i+1\bmod k$ already as an output, then by definition $w_{j}$ changes the output corresponding to the input $j$ from $i$ to $i+1\bmod k$. So there is an occurrence of $i+1\bmod k$ within $(P,[1,n])$. Since $(w_{j}\circ\cdots\circ w_{1})(f)$ cannot have any output more than once, this occurrence is unique.
@@
6 2 1
3 4 2
5 6 3
1 2 4
3 5 6
4 1 2
5 3 4
6 5 1
[…]
4 1 5
6 2 1
3 4 2
5 6 3
1 2 4
3 5 6
4 1 2
5 3 4
6 5 1
[…]
6 2 1
3 4 2
5 6 3
1 2 4
3 5 6
4 1 2
5 3 4
6 5 1
2 6 3
[…]
3 4 2
5 6 3
1 2 4
3 5 6
4 1 2
5 3 4
6 5 1
2 6 3
[…]
3 4 2
5 6 3
1 2 4
3 5 6
4 1 2
5 3 4
6 5 1
2 6 3
4 1 5
6 2 1
3 4 2
5 6 3
1 2 4
3 5 6
4 1 2
5 3 4
6 5 1
2 6 3
4 1 5
Figure 2. We demonstrate, in six panels, how to place a $[6]$-chunk on this $\mathbf{w}$-orbit on $\mathcal{F}=\operatorname{Inj}_{1}(3,6)$. In panel $i$ from left to right, the positions highlighted in green are the first $i$ positions in the chunk. The circled entries (in the first five panels) are the next $n=3$ positions after the one containing the $i$. We choose the unique $i+1$ among the circled entries to be in the chunk.
@@
The idea of the proof is to partition any given orbit into $[k]$-chunks that contain every number $1,2,\dots,k$ exactly once. Within any orbit $\mathcal{O}$, we will assume without loss of generality that 1 appears as an output at least as often as any other number $2,\dots,k$. This is because if $i>1$ appears as an output more times than $1$, then we can renumber $i,i+1,\dots,k,1,\dots,i-1$ as $1,2,\dots,k$, and the outputs remain the same relative to each other (mod $k$).
Choose a 1 in the orbit board and call this position $P_{1}$. Then by Lemma 2.15(1), there exists a unique occurrence of 2 within $(P_{1},[1,n])$; place this 2 in the same chunk as 1 and call its position $P_{2}$. For every $i$ in a chunk in position $P_{i}$, choose the $i+1$ within $(P_{i},[1,n])$ and call its position $P_{i+1}$. Continue this until the chunk contains $1,2,\dots,k$. Refer to Figure 2 for an example of this process. In each step shown, the circled entries are the next $n=3$ positions after the position placed in the chunk.
To start a new chunk, we choose a 1-entry that is not already part of a chunk, and continue the same process. If $Q$ is the position of the $i+1$ in a chunk, the $i$ in the same chunk clearly must be in $Q([-n,-1])$. By Lemma 2.15(2), there is only one $i$-entry that can be in the same chunk as the given $(i+1)$-entry. Thus, our $[k]$-chunks are disjoint.
Once every 1 in the orbit board is part of a completed $[k]$-chunk, there are no more entries not already in a chunk, since we assumed 1 appears as an output at least as often as any other number. The chunking process shows that $1,2,\dots,k$ appear as outputs the same number of times in any orbit. ∎
@@
Now we prove Theorem 2.11 for the case $\mathcal{F}=\operatorname{Inj}_{m}(n,k)$. We use a similar technique as for the $\operatorname{Inj}_{1}(n,k)$ case. We will again partition orbit boards into $[k]$-chunks, where each chunk contains $1,2,\dots,k$ and the instance of $i+1$ within a chunk is at most $n$ positions after the instance of $i$ (in the reading order). Unlike in the $\operatorname{Inj}_{1}(n,k)$ case, there is no longer a unique way of partitioning the orbit into chunks; see Figure 3 for two different ways to partition the orbit of $\operatorname{Inj}_{2}(4,4)$ containing 1441 into $[4]$-chunks. However, all that matters to prove Theorem 2.11 is the existence of a partitioning into $[k]$-chunks. Unfortunately the proof becomes more complicated due to lack of uniqueness of the partition into chunks.
@@
    There are at most $m$ occurrences of $i+1\bmod k$ within $(P,[1,n])$.
@@
    If the position $(P,n)$ directly below $P$ does not contain $i+1\bmod k$, then there are exactly $m$ occurrences of $i+1\bmod k$ within $(P,[1,n-1])$.
@@
    There are at most $m$ occurrences of $i-1\bmod k$ within $(P,[-n,-1])$.
@@
    If the position $(P,-n)$ directly above $P$ does not contain $i-1\bmod k$, then there are exactly $m$ occurrences of $i-1\bmod k$ within $(P,[-(n-1),-1])$.
@@
Suppose position $P$ is in column $j$. Then $f(j)=i$ for some $f\in\mathcal{O}$. By Remark 2.16, there cannot be more than $m$ occurrences of $i+1\bmod k$, proving (1).
If $(w_{j-1}\circ\cdots\circ w_{1})(f)$ does not have $i+1\bmod k$ already $m$ times as an output, then by definition $w_{j}$ changes the output corresponding to the input $j$ from $i$ to $i+1\bmod k$. This proves (2).
@@
By the same relabeling argument from the injections proof, we will assume without loss of generality that within any given orbit $\mathcal{O}$, 1 appears as an output at least as often as any other number $2,\dots,k$.
@@
To start a new chunk, we choose a 1-entry that is not already part of a chunk, and wish to continue the same process. However, unlike in the case of injections, there is not necessarily a unique occurrence of $i+1$ within the next $n$ positions after an $i$, nor a unique occurrence of $i-1$ within the previous $n$ positions before an $i$. When placing an $(i+1)$-entry in the same chunk as $i$ in position $P$, we want to choose $i+1$ that is not already part of a chunk. When $(P,[1,n])$ contains such an $(i+1)$-entry, we choose one of them. However, such an $(i+1)$-entry may not exist depending on how we chose earlier chunks.
See the left side of Figure 4 for an example of this problem. Let $P$ be the position of the 2 in the purple chunk. Then the only 3 in $(P,[1,n])$ is already part of the brown chunk. In this case, we reassign the part of the brown chunk starting with 3 to be in the purple chunk, and then continue where the purple chunk is complete and we now attempt to complete the brown chunk.
In general, suppose for a given $i$ in position $P$, all $(i+1)$-entries within $(P,[1,n])$ are already part of chunks. Then we claim that at least one of these $(i+1)$-entries is in the same chunk as an $i$-entry in $(P,[1,n-1])$. To explain this we consider two cases.

### v4

@@ Abstract.
This paper analyzes a certain action called “whirling” that can be defined on any family of functions between two finite sets equipped with a linear (or cyclic) ordering. Many maps of interest in dynamical algebraic combinatorics, such as rowmotion of order ideals, can be represented as a composition of “toggling” involutions, each of which modifies its object only locally. Similarly whirling is made up of locally-acting whirling maps which directly generalize toggles, but cycle through more than two possible outputs. In this first paper to focus on whirling, we consider it as a map on subfamilies of functions between finite sets.
@@
Keywords: dynamical algebraic combinatorics, homomesy, injection, noncrossing partition, RG-word, set partition, surjection, toggles, toggle group, whirling
@@ 1. Introduction
In this paper, we explore the orbits associated with an action that can be defined on any family $\mathcal{F}$ of functions $f$ from one finite set $S$ to another finite set $T$, where $S$ comes with a linear ordering and $T$ comes with a cyclic ordering. Without loss of generality we may take $S=[n]:=\{1,2,\dots,n\}$ with its usual ordering and $T=[k]$ with the mod $k$ cyclic ordering. The action is generated by an operation $\mathbf{w}:\mathcal{F}\rightarrow\mathcal{F}$ (“the whirling map”) that is in turn defined as the composition of simpler maps $w_{i}:\mathcal{F}\rightarrow\mathcal{F}$ (“whirling at ii”, with $i$ in $[n]$) that repeatedly add $1\bmod k$ to the value of $f$ at $i$ until we get a function in $\mathcal{F}$. Our main results are examples of the homomesy [^1] Greek for “same middle” phenomenon, introduced by the second and third authors [PR15] and defined as follows.
@@
The homomesy phenomenon is quite widespread throughout combinatorial dynamical systems. See [Rob16] for a survey article. Since it is a new area of research over the past roughly 15 years, there is still much to learn about it, including the best techniques for proving homomesy, what types of actions and statistics commonly yield homomesy, and what consequences can arise from homomesy.
Early on, almost all proven homomesy results were for actions whose order was straightforwardly computable in general and relatively small compared to the size of the ground set $\mathcal{S}$. So it is rarer and notable to find homomesies for actions with unpredictable orbit sizes, as is the case with the maps in this paper, which also generally have unknown order. Other such examples can be found in [EFG+16, ELM+24]. In some cases, a statistic which is always an integer has a non-integer average across every orbit, which means all of the orbit sizes must be multiples of the constant average’s denominator. This consequence of homomesy appears in Corollary 2.21.
The whirl maps $w_{i}$ are a generalization of the toggle maps that are heavily studied in dynamical algebraic combinatorics. Toggles were first introduced as a way of writing the rowmotion map as a composition of simpler operations on order ideals [CF95]. Striker later generalized them to a much wider variety of settings [Str18]. There have been many interesting homomesy and periodicity phenomena discovered for actions defined in terms of toggles; see e.g., [PR15, Rob16, Had21, EFG+16, JR18, ELM+24].
In recent work, Plante and the third author consider a whirling action on the set of $k$-bounded $P$-partitions of a poset $P$. Following in the footsteps of Haddadan [Had21], they prove this to be in equivariant bijection with rowmotion acting on order ideals of the poset $P\times[k]$ for any finite poset $P$. They then leverage this to obtain periodicity and homomesy results for rowmotion acting on the poset ${\sf V}\times[k]$, which has surprisingly good dynamical properties [PR24]. A key point here is that beyond the intrinsic interest of whirling as a generalization of toggling, its study can facilitate our understanding of rowmotion questions that arise independently. A related construction is discussed in [DTW03]; we have borrowed the term “whirling” from the paper’s notion of “whirling tours”.
In Section 2 we consider the whirling action on injections and on surjections from a finite set $[n]$ to another finite set $[k]$. In both cases, we prove that within each orbit, any two elements $1,2,\dots,k$ of the codomain appear as outputs of functions the same number of times. We also prove this for a family of functions that naturally generalizes injections and conjecture it for a similar family that generalizes surjections. Then in Section 3, we explore the whirling action on various families of restricted growth words, which are functions that encode set partitions.
@@ 1.1. Acknowledgements
We have benefitted from useful discussions with Yue Cai, David Einstein, Max Glick, Shahrzad Haddadan, Sam Hopkins, Thomas McConville, Matthew Plante, Elizabeth Sheridan Rossi, Richard Stanley, Jessica Striker, and Nathan Williams. Computations leading to the initial conjectures were done in Sage [S+16]. We also thank two anonymous reviewers who have provided helpful feedback.
@@ 2. Whirling on m-injections and m-surjections between finite sets
In this section we first define the discrete functions $m$-injections and $m$-surjections and the whirling action on them, along with the basic tools (e.g., orbit boards) that we use to understand the dynamics. The next few sections detail the lemmas and proofs of homomesy. Then we discuss a consequence of the homomesy for orbit sizes of whirling, for which we have no other proof. We finish with a lifting of the dynamics to the piecewise-linear setting, where the homomesy conjecturally appears to continue to hold.
@@ 2.1. The whirling map on m-injections and m-surjections
A function $f:S\rightarrow T$ is $m$-injective if every element $t\in T$ appears as an output of $f$ at most $m$ times, i.e., $f^{-1}(t)\leq m$ for every $t\in T$. An $m$-injective function is also called an $m$-injection. Injections are the same as 1-injections.
@@
A function $f:S\rightarrow T$ is $m$-surjective if every element $t\in T$ appears as an output of $f$ at least $m$ times, i.e., $f^{-1}(t)\geq m$ for every $t\in T$. An $m$-surjective function is also called an $m$-surjection. Surjections are the same as 1-surjections.
@@
For $f\in\mathcal{F}\subseteq[k]^{[n]}$ define a map $w_{i}:\mathcal{F}\rightarrow\mathcal{F}$, called whirling at index $i$, as follows: repeatedly add 1 (mod $k$) to the value of $f(i)$ until we get a function in $\mathcal{F}$. The new function is $w_{i}(f)$.
Note that we represent our residues (mod $k$) within $[k]$ rather than $\{0,1,2,\dotsc,k-1\}$. The outputs of a function in $\mathcal{F}$ are always to be considered mod $k$, so we will not continue to write “mod kk.”
@@
Let $\mathcal{F}=\operatorname{Inj}_{2}(6,4)$ and $f=422343$. To compute $w_{3}(f)$, we first add 1 to $f(3)=2$ and get the function $423343$, which is not 2-injective since 3 appears as an output three times. So, we add 1 again to the third position and get $424343$, which is also not 2-injective. Adding 1 once more gives a 2-injective function $421343=w_{3}(f)$.
@@
$2753\stackrel{{\scriptstyle w_{1}}}{{\longmapsto}}4753\stackrel{{\scriptstyle w_{2}}}{{\longmapsto}}4153\stackrel{{\scriptstyle w_{3}}}{{\longmapsto}}4163\stackrel{{\scriptstyle w_{4}}}{{\longmapsto}}4165$
@@
On any family $\mathcal{F}$, $w_{i}$ and $\mathbf{w}$ are invertible. Given $f\in\mathcal{F}$, we get $w_{i}^{-1}(f)$ by repeatedly subtracting 1 from the value of $f(i)$ until we get a function in $\mathcal{F}$.
@@
Conjecture 2.11 has been verified by a computer for all triples $(m,n,k)$ where $n\leq 10$ (for which there are finitely many relevant triples since $\operatorname{Sur}_{m}(n,k)=\O$ if $mk>n$).
2.2. Orbit board terminology and notation
In this subsection, we introduce terminology and notation that applies to whirling on any family of functions $\mathcal{F}$ from $[n]$ to $[k]$ that we will use on various homomesy proofs. We use a generalization of the tuple boards of Haddadan; see[Had21].
Let $\mathcal{O}$ be an orbit under the action of $\mathbf{w}$ on $\mathcal{F}$. We draw a board for the orbit $\mathcal{O}$ by placing some $f\in\mathcal{O}$ on the top row. The function in row $i+1$ is $\mathbf{w}^{i}(f)$ for $i\in[0,\ell(\mathcal{O})-1]$, where $\ell(\mathcal{O})$ is the length of $\mathcal{O}$. For example, in Figure 1, we show a board for an orbit on $\mathcal{F}=\operatorname{Inj}_{1}(3,6)$. From this board, we see that if we let $f=621$, then $\mathbf{w}(f)=342$, $\mathbf{w}^{2}(f)=563$, and so on, and that $\mathbf{w}^{10}(f)=f$.
Notice that within an orbit board, if $f$ is a given row, then the “partially whirled element” $(w_{i}\circ\cdots\circ w_{1})(f)$ is given by the first $i$ numbers on the row below $f$ and the last nin-i numbers of $f$. (We say the row “below” the bottom row is the top row, as we consider the orbit board to be cylindrical.) For instance, in the first two rows of Figure 1, $f=621$, $w_{1}(f)=321$, and $(w_{2}\circ w_{1})(f)=341$.
@@ 2.2. Orbit board terminology and notation
Figure 1. The orbit under the action of $\mathbf{w}$ on $\operatorname{Inj}_{1}(3,6)$ containing $f=621$.
We use the term reading the orbit board to refer to this action where we start at a certain position $P$ of the board, and continue to the right until we reach the end of the row, and then go to the leftmost position of the row below and continue. When we refer to a certain position being $x$ positions “before” or “after” the position $P$, or say the “previous” or “next” position, we always mean in the reading order.
@@
For a given position $P$ in an orbit board, let $(P,h)$ denote the position $h$ places after $P$ in the reading order (or h-h places before $P$ if $h<0$). Let $(P,[a,b])$ denote the (ba+1)(b-a+1)-tuple $\left((P,a),(P,a+1),\dots,(P,b-1),(P,b)\right)$.
@@
Consider the following orbit on $\operatorname{Inj}_{1}(4,5)$. Let $P$ be the position in the second row and second column, shown below surrounded by a black rectangle. Then $(P,[1,4])$ consists of the four positions circled in red. Also $(P,[0,4])$ is $(P,[1,4])$ together with $P$, while (P,[1,2])(P,[-1,2]) is the second row. As the orbit is cylindrical, the bottom right corner is both $(P,14)$ and (P,6)(P,-6). Note that $P$ refers only to the position, not the value in that position. So $P\not=(P,5)$ since they are different positions, even though both contain the value 3. Similarly, we will never write, e.g., $P=3$.
@@
Note that if $P$ is in row $i$ and column $j$, then $(P,[1,n])$ always consists of all positions to the right of $P$ in row $i$, followed with the leftmost $j$ positions of row $i+1$. Also note that P([n,1])P([-n,-1]) consists of the rightmost nj+1n-j+1 positions of row i1i-1 followed by all j1j-1 positions left of $P$ in row $i$.
Remark 2.14.
2.3. The proof of Theorem 2.10 for injections
In this subsection, we prove Theorem 2.10 when $\mathcal{F}=\operatorname{Inj}_{1}(n,k)$. Even though this is a special case of $m$-injections, the proof here is much simpler.
@@
    There is exactly one occurrence of $i+1$ within $(P,[1,n])$.
@@
    There is exactly one occurrence of i1i-1 within (P,[n,1])(P,[-n,-1]).
@@
To prove (1), suppose position $P$ is in column $j$. So $f(j)=i$ for some $f\in\mathcal{O}$. Then $(P,[1,n])$ contains the multiset of outputs of $(w_{j}\circ\cdots\circ w_{1})(f)$ in some order. If $(w_{j-1}\circ\cdots\circ w_{1})(f)$ does not have $i+1$ already as an output, then by definition $w_{j}$ changes the output corresponding to the input $j$ from $i$ to $i+1$. So there is an occurrence of $i+1$ within $(P,[1,n])$. Since $(w_{j}\circ\cdots\circ w_{1})(f)$ cannot have any output more than once, this occurrence is unique.
@@
By Lemma 2.15, we have a bijection between $i$-entries and $(i+1)$-entries for any $i\in[k-1]$, leading immediately to the equidistribution. ∎
@@
Now we prove Theorem 2.10 for the case $\mathcal{F}=\operatorname{Inj}_{m}(n,k)$. We will partition orbit boards into $[k]$-chunks, where each chunk contains $1,2,\dots,k$ and the instance of $i+1$ within a chunk is at most $n$ positions after the instance of $i$ (in the reading order). Note that there is not necessarily a unique way of partitioning the orbit into chunks; see Figure 2 for two different ways to partition the orbit of $\operatorname{Inj}_{2}(4,4)$ containing 1441 into $[4]$-chunks. However, to prove Theorem 2.10 we simply need to show the existence of at least one partitioning into $[k]$-chunks.
@@
    There are at most $m$ occurrences of $i+1$ within $(P,[1,n])$.
@@
    If the position $(P,n)$ directly below $P$ does not contain $i+1$, then there are exactly $m$ occurrences of $i+1$ within (P,[1,n1])(P,[1,n-1]).
@@
    There are at most $m$ occurrences of i1i-1 within (P,[n,1])(P,[-n,-1]).
@@
    If the position (P,n)(P,-n) directly above $P$ does not contain i1i-1, then there are exactly $m$ occurrences of i1i-1 within (P,[(n1),1])(P,[-(n-1),-1]).
@@
Suppose position $P$ is in column $j$. Then $f(j)=i$ for some $f\in\mathcal{O}$. By Remark 2.16, there cannot be more than $m$ occurrences of $i+1$, proving (1).
If $(w_{j-1}\circ\cdots\circ w_{1})(f)$ does not have $i+1$ already $m$ times as an output, then by definition $w_{j}$ changes the output corresponding to the input $j$ from $i$ to $i+1$. This proves (2).
@@
Within any orbit $\mathcal{O}$, we will assume without loss of generality that 1 appears as an output at least as often as any other number $2,\dots,k$. This is because if $i>1$ were to appear as an output more times than $1$, then we could renumber $i,i+1,\dots,k,1,\dots,i-1$ as $1,2,\dots,k$, and the outputs would remain the same relative to each other (mod $k$).
@@
To start a new chunk, we choose a 1-entry that is not already part of a chunk, and wish to continue the same process. However, there is not necessarily a unique occurrence of $i+1$ within the next $n$ positions after an $i$, nor a unique occurrence of i1i-1 within the previous $n$ positions before an $i$. When placing an $(i+1)$-entry in the same chunk as $i$ in position $P$, we want to choose $i+1$ that is not already part of a chunk. When $(P,[1,n])$ contains such an $(i+1)$-entry, we choose one of them. However, such an $(i+1)$-entry may not exist depending on how we chose earlier chunks.
See the left side of Figure 3 for an example of this problem. Let $P$ be the position of the 2 in the purple chunk. Then the only 3 in $(P,[1,n])$ is already part of the brown chunk. In this case, we reassign the part of the brown chunk starting with 3 to be in the purple chunk. So the purple chunk is now complete but the brown chunk is incomplete, having just the values 1 and 2. We now need to complete the brown chunk.
@@
Claim: Suppose for a given $i$ in position $P$, all $(i+1)$-entries within $(P,[1,n])$ are already part of chunks. Then at least one of these $(i+1)$-entries is in the same chunk as an $i$-entry in (P,[1,n1])(P,[1,n-1]).
To explain why the claim is true, we consider two cases